How can I solve the equations: 3x+y=3, 7x+2y=1
Answer by mukhopadhyay(490)

3x+y=3
=>y=3(1-x)
7x+2y=1
=>7x+2*3(1-x)=1
=>x=-5;
y=3(1-x) => y=18
Answer: x=-5 and y=18